OpenStudy (anonymous):

Find the value of b using the given information A-5.4=8.3 and b=4a

OpenStudy (jhannybean):

Is \(A\) the same thing as \(a\)?

OpenStudy (anonymous):

so b=4a if we rearrange it we will get

OpenStudy (anonymous):

\[a=\frac{ b }{ 4 }\]

OpenStudy (anonymous):

now \[a-5.4=8.3\]

OpenStudy (anonymous):

plug in the value of a in this equation

OpenStudy (anonymous):

\[\frac{ b }{ 4 }-5.4=8.3\]

OpenStudy (anonymous):

now solve for b , let me know if u need help with that too :)

OpenStudy (anonymous):

Need help

OpenStudy (anonymous):

ok so \[\frac{ b }{ 4}=8.3+5.4\]

OpenStudy (anonymous):

\[\frac{ b }{ 4 }=13.7\]

OpenStudy (anonymous):

\[b=13.7 \times 4\]

OpenStudy (anonymous):

\[b=54.8\]